find the point of intersction forthe line below using elimation 3x+4y=-2, 2x+5y=1
Answer by checkley71(8403)

multiply the first equation by 2 & the second by -3 & then add thus
6x+8y=-4
-6x-15y=-3
--------------------------
-7y=-7
y=-7/-7
y=1 solution. now substitute 1 for y in either equation & solve for x
3x+4*1=-2
3x+4=-2
3x=-2-4
3x=-6
x=-6/3
x=-2 solution.
proof
2*-2+5*1=1
-4+5=1
1=1